A baker is using a cookies recipe that callc for 2 1/4 cups of flour to yield 36 cookies. how much fliur will the baker need to make 90 cookes using the same recipe?
Answer #1

one cup of flour (wheat)= 120 gm

hence 2 and 1/4 cups= 240+30= 270 gm

Given that for 36 cookies required flour is= 2 and 1/4 cup flour= 270 gm

hor 1 cookies = 270/36

Hence for 90 cookies required flour = (270/36) X90

=675 gm of flour

or 675/120=5.625 cups